Maximize x+y+z subject to
x+y is x+z is y+z is Also know that x,y,z is >or= 0

Answer by stanbon(75887)

Maximize x+y+z subject to
x+y is x+z is y+z is Also know that x,y,z is >or= 0
---------------------
x + y + z = (x+y)+z = (x+z)+z = x+2z
-----
x+y+z = (x+y)+z = y+z + z = y+2z
----
So, x+2z = y+2z
Therefore x = y
-------------------
Cheers,
Stan H.
-----------------